# IEEEexample:BSTcontrol

### v3

@@ 3.2 Interpolation Consistency
Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timpoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations. By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
@@ 3.4 Unlimited Sampled Triplets
As illustrated in Figure˜3 (Right), the proposed inter-/extra-polation constraints require observed triplets of timepoints, limiting learning when few discrete scans are available. To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals. Specifically, by sampling intermediate timepoints $\hat{t}$ from $\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints. This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
@@ 4 Data and Experiments
We train and evaluate TimeFlow on the public Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [4], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both healthy subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders that exhibit specific patterns of atrophy development in the brain. We employ the longitudinal processing from FreeSurfer v7.2 [49]. First, all scans are processed individually, including skull stripping, normalization. Subsequently, the image sequence is rigidly aligned via a groupwise rigid registration [50]. FreeSurfer aggregates information across time points and refines the pre-processed outputs. To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm). Then, the dataset is divided into training, validation, and testing sets, ensuring a balance in diagnosis, age, sex, and the number of visits. It results in 74, 30, 30 subjects for each split, respectively, with the number of time points per subject ranging from 2 to 12. All scans have isotropic 1mm spacing and are cropped to a uniform size of $160\times 160\times 192$ voxels. The intensity values are scaled to the range [0,1], from percentile [0,99.9] without clipping. Additionally, 11 subjects from the OASIS dataset [51] serve for zero-shot evaluation.
@@
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per subject). We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
@@
To evaluate the registration performance, we compare TimeFlow with the following methods:
  • NODER [17, 16] takes the complete image sequence as the model input and is thus aware of every time point. It carries out subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [52] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [25] and the diffeomorphic variant (VXMdiff[53]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The number of integration steps of VXMdiff is also set to 7.
  • Sequence-based VoxelMorph (Seq-VXM) [15]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate time point. Given that Seq-VXM performs multiple pairwise registrations by taking the full sequence as model input, we treat Seq-VXM as the performance upper bound for the retrospective study (refer to Tab. 1).
All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss. The loss function weighting factors in NODER follow their original suggestion. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [54, 55] with weight $1.5,0.5,1.0$.
@@ 5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans. TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a surrogate metric to quantify the biological aging progression.
@@
A significant advantage of TimeFlow, as described in Section˜3, is its ability to extrapolate deformation fields to future timepoints beyond the observed interval. The quantitative performance of TimeFlow’s extrapolation forecasting is reported in Table˜1. The metrics are computed with respect to the ground truth future scans. Results clearly demonstrate that TimeFlow consistently achieves superior extrapolation accuracy compared to baseline methods, yielding the lowest MAE and highest PSNR (Figure˜8), even when evaluated in a zero-shot manner on the unseen OASIS dataset. Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison. Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images. Moreover, Figure˜5 demonstrates extrapolation at a large $t$ ($t=2,4,6$) where ground truth is not available. In these cases, NODER produces unrealistic and anatomically implausible deformations, resulting in significant errors in future image prediction. Similarly, VXM/diff, which applies linear scaling to displacement/velocity fields, generates unsmooth deformations (indicated by high NDV) and dissolves the brain, causing the anatomical details to fade. In contrast, TimeFlow/diff consistently generates smooth and coherent extrapolated deformation fields, quantitatively supported by the lowest MAE and minimal NDV values in Table˜1.
@@
As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis. We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans. Specifically, MRI triplets are sampled from validation and test subsets, where the intervals between the first and second visits match closely the intervals between the second and third visits. The diagnosis at the third visit labels each triplet. Using the first two scans as input, TimeFlow extrapolates forward to the third, future scan. By comparing predictions against actual observed future, we compute the extrapolated time parameter $t_{\textrm{ext}}$, and divide it by the normalized chronological intervals (in years) to yield a brain aging rate. An aging rate above 1 indicates accelerated pathological changes, as expected in patients with neurodegenerative conditions. Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1. The analysis validates that our annotation-free surrogate brain aging rate metric is able to separate disease groups.
@@ 5.2 Retrospective studies: Matching the deformation trajectory
We further evaluate TimeFlow’s interpolation capability, assessing its effectiveness in modeling temporally continuous and anatomically plausible deformation trajectories connecting observed MRI visits. In this setting, TimeFlow uses only the baseline visit $I_{0}$ and the final visit $I_{L}$ as inputs, predicting deformation fields $\phi_{t_{k}}$ for intermediate images $I_{k}$, where $0<k<L$. Quantitative comparisons with baseline methods for interpolation are presented in Table˜1. As discussed in baselines, Seq-VXM, performing explicit pairwise registrations for every available timepoint, represents a performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER on the ADNI dataset, which is also aware of the entire sequence and parametrized by Neuralordinary differential equation (ODE). TimeFlow achieves comparable performance to NODER in zero-shot evaluation on the OASIS dataset, where NODER performs subject-specific optimization on the full image sequence. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization. Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments. Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a surrogate brain progression state marker. We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment). The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
@@ 5.3 Ablation Study
As described in Section˜3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit smoothness regularization. To assess the effectiveness of our proposed inter-/extra-polation consistency constraints, we perform detailed ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric. As shown in table˜2, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation. These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.

### v4

# Temporal Conditioning for Longitudinal Brain MRI Registration and Aging Analysis

@@ 3.2 Interpolation Consistency
Here, both warped images $I_{0}\circ\phi_{\hat{t}}^{0\rightarrow L}$ and $I_{L}\circ\phi_{\hat{t}-1}^{0\rightarrow L}$ serve as a synthetic version of intermediate timepoint $I_{\hat{t}}$, eliminating reliance on actual intermediate observations. By freely sampling $\hat{t}$, TimeFlow effectively learns densely and continuously across the entire temporal axis.
@@ 3.4 Unlimited Sampled Triplets
As illustrated in Figure˜3 (Right), the proposed inter-/extra-polation constraints require observed triplets of timepoints, limiting learning when few discrete scans are available. To address this, we use “synthetic” triplets dynamically sampled from continuous time intervals. Specifically, by sampling intermediate timepoints from uniform distribution $\hat{t}\sim\mathcal{U}(0,1)$ during training, we can generate infinite triplets from a pair of source and target images, freeing TimeFlow from the discrete, fixed timepoints. This continuous, densely sampled temporal supervision significantly enhances the network’s capacity to generalize and accurately model brain anatomy changes at any arbitrary timepoint, including unseen future scenarios, while using only two scans as inputs.
@@ 4 Data and Experiments
We train and evaluate TimeFlow on the publicly available Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [21], which comprises longitudinal T1-weighted Magnetic Resonance (MR) scans from both cognitively normal subjects and patients diagnosed with Alzheimer’s Disease (AD) or Mild Cognitive Impairment (MCI), i.e., neurodegenerative disorders characterized by progressive patterns of brain atrophy. All scans undergo standardized pre-processing using the longitudinal pipeline in FreeSurfer v7.2 [13]. Each scan is first processed individually for skull stripping, normalization. Subsequently, all scans from the same subject are rigidly aligned via a groupwise rigid registration [41]. FreeSurfer aggregates information across timepoints and refines the pre-processed outputs. To ensure meaningful longitudinal analysis, we select 134 subjects exhibiting substantial structural changes (median brain deformation >0.3>0.3mm). The dataset is partitioned into training (74 subjects), validation (30 subjects), and testing (30 subjects) sets at the subject level prior to any processing, ensuring complete subject disjointness across splits and preventing information leakage. The splits are constructed to maintain balance across diagnosis, age, sex, and number of available visits. Each subject contributes to 2 to 12 longitudinal scans. All images are resampled to isotropic 1mm spacing and cropped to a uniform size of $160\times 160\times 192$ voxels. Intensity values are normalized to the range [0,1] based on the [0, 99.9] percentile without clipping. Additionally, 11 subjects from the OASIS dataset [24] are used for zero-shot evaluation to assess TimeFlow’s generalization to data from different scanners and imaging protocols.
@@
Our model is implemented using Pytorch 2.1.2. and training/test is carried out on one NVIDIA A100 40G with a batch size of $2$ (sample $\hat{t}$ twice per image pair). We set the dimension of time embedding to $16$, the feature channels in the U-Net encoder and decoder to $[32,32,48,48,96]$. In addition to TimeFlow, which outputs the time-dependent deformation field directly, we have the diffeomorphic variant TimeFlowdiff, which utilizes the scaling and squaring layer with $7$ integration steps to integrate the temporal stationary velocity field. The weights of both interpolation and extrapolation similarity are set to $1.0$. The weights of interpolation and extrapolation flow consistency are set to $2,0.03$ for TimeFlow and $1.25,0.025$ for TimeFlowdiff.
@@
To evaluate TimeFlow’s registration performance, we compare against the following methods:
  • NODER [54, 3] takes the complete image sequence as input and is thus aware of every timepoint. It performs subject-specific optimization and models the time-varying velocity field with a neural network and leverages NeuralODE [8] to compute integration and adjoint gradients. It fits the subject-specific temporal deformation trajectory by maximizing the similarity with each frame.
  • VoxelMorph (VXM) [4] and the diffeomorphic variant (VXMdiff[10]. Since both are pairwise registration methods, the interpolation and extrapolation prediction is carried out based on the linear scaling of the displacement or velocity field from the pairwise registration. The number of integration steps of VXMdiff is set to 7.
  • SyN (ANTs) [1]. We apply the symmetric diffeomorphic image pairwise registration algorithm with normalized cross-correlation (radius $r=3$). For interpolation and extrapolation, displacement fields are linearly scaled, analogous to VXM.
  • Sequence-based VoxelMorph (Seq-VXM) [53]. Seq-VXM processes the whole sequence and registers the baseline image to each intermediate timepoint through multiple pairwise registration. Since Seq-VXM uses the full image sequence as input, we treat it as the performance upper bound for the retrospective interpolation task (refer to Tab. 1).
All methods use Local Normalized Cross Correlation (LNCC) with radius $r=9$ (NODER) or $r=4$ (others) as image similarity loss. The loss function weighting factors for NODER follow the original implementation. Seq-VXM, VXM and VXMdiff compute the SOTA smoothness regularizer GradICON [48, 47] with weight $1.5,0.5,1.0$.
@@
Table 2: Dice Score on ADNI dataset. For each method, we report the mean Dice scores and standard deviation across all subjects and anatomical regions. Higher values indicate better alignment.
Method Extrapolation Interpolation
Seq-VXM - 81.17 $\pm$ 1.85
SyN 72.85 $\pm$ 6.03 80.32 $\pm$ 2.76
NODER 75.43 $\pm$ 4.43 80.43 $\pm$ 1.88
VXM 75.78 $\pm$ 4.48 81.15 $\pm$ 3.14
VXMdiff 75.35 $\pm$ 4.50 81.11 $\pm$ 2.75
TimeFlow 76.58 $\pm$ 3.90 81.19 $\pm$ 2.88
TimeFlowdiff 76.95 $\pm$ 3.80 81.10 $\pm$ 2.65
@@ 5.1 Prospective studies: Predicting the future
We first evaluate TimeFlow’s performance in prospective scenarios, specifically, its ability to extrapolate deformation fields to predict future brain scans. TimeFlow supports two forms of prospective analysis: (1) Given two available scans and a future time variable $t>1$, the model directly predicts the corresponding deformation fields and resulting future MRI images; (2) Given an observed future scan, TimeFlow can infer the corresponding $\phi_{t}$ and use the estimated time variable $t$ as a quantitative metric to quantify the biological aging progression.
@@
A significant advantage of TimeFlow, as described in Section˜3, is its ability to extrapolate deformation fields to future timepoints beyond the observed interval. The quantitative performance of TimeFlow’s extrapolation forecasting is reported in Table˜1. The metrics are computed with respect to the ground truth future scans. Results clearly demonstrate that TimeFlow consistently achieves superior extrapolation accuracy compared to baseline methods, yielding the lowest MAE and highest PSNR (Figure˜8), even when evaluated in a zero-shot manner on the unseen OASIS dataset. Note that Seq-VXM cannot carry out future predictions; thus, it is excluded from this comparison. In addition, the Dice score metric reported in Tables˜2 and 10 further confirms, in agreement with the intensity-based metrics, that TimeFlow achieves better anatomical region alignment. Qualitative extrapolation results are visualized in Figure˜6 (Year 8) and Figure˜4, illustrating TimeFlow’s accuracy by comparing to the ground truth images. Moreover, Figure˜5 demonstrates extrapolation at a large $t$ ($t=2,4,6$) where ground truth is not available. In these cases, NODER produces unrealistic and anatomically implausible deformations, resulting in significant errors in future image prediction. Similarly, VXM/diff, which applies linear scaling to displacement/velocity fields, generates unsmooth deformations (indicated by high NDV) and dissolves the brain, causing the anatomical details to fade. In contrast, TimeFlow/diff consistently generates smooth and coherent extrapolated deformation fields, quantitatively supported by the lowest MAE and minimal NDV values in Table˜1.
@@
As emphasized in Section˜1, the extrapolation capability of TimeFlow enables annotation-free disease progression analysis (i.e., requiring no anatomical segmentation or surface reconstructions for quantification). We compare the brain aging rates across diagnostic groups, i.e., Controls (CN), MCI, and Dementia, by predicting the optimal time parameter ($t_{\textrm{ext}}$) corresponding to the observed future MRI scans. Specifically, MRI triplets are sampled from validation and test subsets, where the intervals between the first and second visits match closely the intervals between the second and third visits. The diagnosis at the third visit labels each triplet. Using the first two scans as input, TimeFlow extrapolates forward to the third, future scan. By comparing predictions against actual observed future, we compute the extrapolated time parameter $t_{\textrm{ext}}$, and divide it by the normalized chronological intervals (in years) to yield a brain aging rate. An aging rate above 1 indicates accelerated pathological changes, as expected in patients with neurodegenerative conditions. Indeed, the boxplot of Figure˜7 (Left) clearly shows that Dementia patients exhibit the highest progression rate, followed by MCI subjects, while the control group maintains progression rates around 1. The analysis validates that our annotation-free brain aging rate estimate is able to separate disease groups.
Figure 9: Stability analysis across interpolation and extrapolation regimes. Left: Mean Absolute Error (MAE, in units of $\times 10^{-2}$) of TimeFlowdiff as a function of time $t$. The solid line shows the Gaussian kernel-smoothed mean, and the shaded region represents $\pm 1$ standard deviation across subjects. Right: Non-diffeomorphic Volume (NDV, logarithmic scale) of TimeFlowdiff as a function of time $t$. Hollow dots indicate NDV =0=0 (perfectly diffeomorphic). The vertical dashed line at $t=1$ separates interpolation ($0<t<1$) from extrapolation ($t>1$). The x-axis uses a piecewise linear scale: the left segment spans $t\in[0,1]$ and the right segment spans $t\in[1,10]$.
Figure 10: Region-wise Dice score comparison for extrapolation on ADNI dataset. Each boxplot shows the distribution of Dice scores across subjects for major anatomical structures. TimeFlow variants achieve consistently higher or comparable Dice scores across all regions.
Figure 11: Region-wise Dice score comparison for interpolation on ADNI dataset. Each boxplot shows the distribution of Dice scores across subjects for major anatomical structures. All methods achieve high Dice scores ($80\%$), with TimeFlow variants performing comparably to the upper-bound Seq-VXM.
@@ 5.2 Retrospective studies: Matching the deformation trajectory
We further evaluate TimeFlow’s interpolation capability, assessing its effectiveness in modeling temporally continuous and anatomically plausible deformation trajectories connecting observed MRI visits. In this setting, TimeFlow uses only the baseline visit $I_{0}$ and the final visit $I_{L}$ as inputs, predicting deformation fields $\phi_{t_{k}}$ for intermediate images $I_{k}$, where $0<k<L$. Quantitative comparisons with baseline methods for interpolation are presented in Table˜1. As discussed in baselines, Seq-VXM, performing explicit pairwise registrations for every available timepoint, represents a performance upper bound. Despite using only the baseline and last scan as inputs, TimeFlow outperforms NODER on the ADNI dataset, which is also aware of the entire sequence and parametrized by Neuralordinary differential equation (ODE). TimeFlow achieves comparable performance to NODER in zero-shot evaluation on the OASIS dataset, where NODER performs subject-specific optimization on the full image sequence. Additionally, TimeFlowdiff achieves the least non-diffeomorphic volume without extra spatial smoothness regularization. Region-wise segmentation-based evaluation (Tables˜2 and 11) validates the superior performance of TimeFlow and TimeFlowdiff. Qualitative visualizations in Figure˜6 (Year 1.5) illustrate the accuracy of TimeFlow’s interpolations, substantially reducing alignment errors compared to initial misalignments. Similar to our above prospective analysis, we also perform retrospective, group-wise brain progression analysis using the inferred temporal parameter $t$ of intermediate visits as a quantitative index of brain progression state marker. We extract trajectory segments with the same diagnoses from each subject, and then compute a linear relationship between the estimated brain state ($t$) and chronological age (years since baseline of the segment). The median linear lines of the groups (CN, MCI, Dementia) in Figure˜7 (Right) clearly illustrate distinct progression patterns, with Dementia exhibiting the fastest structural changes, MCI intermediate, and CN the slowest.
5.3 Clinical Validation, Interpretability and Robustness
Correlation with Cognitive Decline
To further validate the clinical relevance of TimeFlow’s learned trajectories, we assess the association between the model’s estimated brain state ($t$) and cognitive performance using Repeated Measures Correlation (rmcorr). This analysis reveals a significant negative correlation (r=0.46r=-0.46, 95% CI: [0.51,0.40][-0.51,-0.40]) between the inferred brain state ($t$) and Mini-Mental State Examination (MMSE) scores, indicating that greater pathological progression is associated with declining cognitive function. Importantly, this correlation emerges from trajectories constructed purely from anatomical features of the input image pairs, without any supervision from cognitive scores or diagnostic labels during training. This finding demonstrates that TimeFlow’s unsupervised deformation modeling effectively captures clinically meaningful, subject-specific disease progression patterns.
Distinction from diagnostic classification
Unlike supervised diagnostic classification models that predict discrete disease labels, TimeFlow operates as an unsupervised and disease-agnostic framework designed to capture continuous spatiotemporal structural changes. The model receives no explicit disease labels during training. Instead, it learns to model anatomical trajectories from image pairs alone. The strong alignment between TimeFlow’s derived progression metrics and diagnostic groups (Figure˜7) emerges naturally from the learned deformation fields, confirming that the framework effectively quantifies pathological deviations without disease-specific priors. This unsupervised nature enables TimeFlow to provide granular, interpretable progression quantification that complements categorical diagnoses. Moreover, by generating predicted future brain images, TimeFlow offers clinicians direct visual intuition: "if the subject continues progressing at the current rate, what will their brain look like in, e.g., 2 years?" This capability bridges quantitative biomarkers with interpretable visual assessment, supporting more informed clinical decision-making.
Stability of extrapolation
To assess TimeFlow’s robustness across both interpolation and long-range extrapolation, we analyze registration accuracy (MAE) and deformation regularity (NDV) as functions of time $t$ in Figure˜9. For interpolation ($0<t<1$), TimeFlow achieves consistently low MAE (below $8\times 10^{-2}$) and diffeomorphic deformations (NDV =0=0). For extrapolation ($t>1$), MAE increases monotonically with prediction distance, reflecting the inherent uncertainty in long-term forecasting. However, for near-to-mid-term predictions ($1<t\leq 5$), the error remains stable below $10\times 10^{-2}$, comparable to the mean extrapolation MAE reported in Table˜1. Crucially, NDV remains negligible ($0.01\%$ of total brain volume) for $t\leq 5$, indicating minimal folding artifacts. These results confirm that TimeFlow generates anatomically plausible deformations and maintains stability for clinically relevant prediction horizons.
@@ 5.4 Ablation Study
As described in Section˜3 and supported by both quantitative and qualitative results, TimeFlow achieves temporally continuous and spatially smooth deformation fields without requiring explicit smoothness regularization. To assess the effectiveness of our proposed inter-/extra-polation consistency constraints, we perform detailed ablation studies. To evaluate the temporal smoothness and continuity of TimeFlow, we compute the derivative $\partial\phi_{t}/\partial t$ using the finite difference method by sampling $t$ at intervals of $\Delta t=0.25$. Similar to the commonly used diffusion regularizer in learning-based registration methods, which computes the Dirichlet energy over spatial gradients, we quantify temporal smoothness as $\|\partial\phi_{t}/\partial t\|$. Specifically, we compute the norm of the temporal displacement derivatives within the brain regions and then take the mean over all brain voxels as the final metric. As shown in table˜3, incorporating interpolation consistency loss significantly improves both temporal and spatial smoothness ($\|\partial\phi_{t}/\partial t\|$ and NDV drastically reduce) without sacrificing interpolation accuracy. Furthermore, the extrapolation similarity and flow consistency losses enable TimeFlow to generate plausible extrapolated deformation fields for future image estimation. These ablation results demonstrate that TimeFlow inherently learns temporally continuous and spatially smooth deformation fields, achieving an optimal balance between registration accuracy and smoothness without explicit smoothness penalties.
5.5 Runtime analysis
We compare computational efficiency across all methods. Optimization-based methods exhibit substantially longer runtimes: SyN requires approximately 5 minutes per image pair on CPU, while NODER performs subject-specific optimization requiring 1-2 minutes per subject on GPU. In contrast, learning-based methods achieve efficient amortized inference. TimeFlow and TimeFlowdiff complete registration in 5-7 seconds per subject, VXM and VXMdiff in 1-2 seconds, and Seq-VXM in 3-5 seconds (all on GPU). Beyond runtime, memory requirements differ significantly for long sequences: NODER and Seq-VXM require over 16 GB GPU memory for sequences exceeding 5 timepoints, whereas TimeFlow operates efficiently with less than 8 GB.